Grade 11Trigonometry

Cot(a).cot(60+a).cot(60-a)=????????????????????????????

Himanshu Dahiya
7 Years agoGrade 11

3 Answers

Arun
7 Years ago
 
Dear Himanshu
 
it can be written as 1/ [ tan A tan (60-A) tan (60+A)] =1/ tan 3A
 
hence = cot 3A
 
Hope it helps
 
In case of any difficulty, please feel free to ask.
Mohammad Hesham
7 Years ago
Cota.Cot(60+a).Cot(60-a) =cosa.cos(60+a)cos(60-a)/sina.sin(60+a)sin(60-a)=cosa{2cos(60+a).cos(60-a)}/sina{2sin(60+a)sin(60-a)} =cosa.(cos2a+cos120)/sina(cos2a-cos120)=cosa.cos2a+cosa.cos120/sina.cos2a-sina.cos120=2cosa.cos2a-cosa/2sina.cos2a+sina=cosa+cos3a-cosa/sin3a-sina+sina=cos3a/sin3a=cot3a
TEJA KRISHNA
6 Years ago
 
Dear Himanshu
It can be written as 1/ [ tan A tan (60-A) tan (60+A)] =1/ tan 3A
 
hence = cot 3A
all the best
hope u understand this